OpenStudy (anonymous):

Please help me come up with a system of equations! The question is: Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? -- I know how to solve a system, I just need help finding out what the system is for this equation

OpenStudy (austinl):

\[x+y=46\] \[0.10x+0.25y=7.00\] I believe that this is correct.

OpenStudy (anonymous):

thank you! I will try it out now!

OpenStudy (anonymous):

From the first equation, I subtracted x from both sides leaving me with y = 46 - x I put that into the second equation with .10x + .25(46 - x) = 7

OpenStudy (anonymous):

I distributed and got .10d + 11.5 - .25d = 7...but I feel like I'm doing something wrong?

OpenStudy (anonymous):

sorry. by d i meant x

OpenStudy (austinl):

I did it through elimination. Perhaps if you try it that way it will work out better.

OpenStudy (anonymous):

did you multiply the top equation by .10? for elimination

OpenStudy (austinl):

I multiplied the top equation by -0.10

OpenStudy (austinl):

Have you gotten anywhere?

OpenStudy (anonymous):

Is 30 dimes and 16 quarters the correct answer?

OpenStudy (austinl):

That is what I got as a final answer!

OpenStudy (anonymous):

great! Thanks so much for all the help! I really appreciate it!

OpenStudy (austinl):

No problem! If you need any more help, feel free to tag me. I love these type of problems!

OpenStudy (anonymous):

well now that you mention it... I do have another problem...

OpenStudy (austinl):

Ok! Algebra is fun to me :)

OpenStudy (anonymous):

I'm about to post it in the open questions, that way if you help again at least i will be able to give you another medal (:

OpenStudy (austinl):

Ok, sounds good :)